OpenStudy (anonymous):

Show that \[\left( 2x+3\sqrt{x} \right)^{2}\] Can be written as \[4x ^{2}+P \sqrt{x ^{3}}+9x\] Where P is a constand to be found

hartnn (hartnn):

either you can foil it, (2x+3 root x) * (2x+3 root x) to get that form, or you can directly apply the formula \((a+b)^2= a^2+2ab+b^2\)

OpenStudy (anonymous):

I did the first way (2x+3 root x) * (2x+3 root x) Which gave me \[4x ^{2}+12x ^{1.5}+9x\] Now I'm not sure what to do

OpenStudy (loser66):

1.5 = 3/2 when putting it under the root you have sqrt ( x^3)

hartnn (hartnn):

thats correct, so your P is 12, right ? the co-efficient of the middle term

hartnn (hartnn):

what exactly was your doubt ? because you were done :P

hartnn (hartnn):

\(\sqrt{x^3}= (x^3)^{(1/2)}=x^{(3/2)}=x^{1.5}\)

OpenStudy (anonymous):

Well that wasn't so bad, I was expecting to have to do something else to the equation. Thanks again hartnn.

hartnn (hartnn):

oh,ok. welcome ^_^